OpenStudy (anonymous):

How to solve 7x-1/2(8x+4)=6 step by step

OpenStudy (anonymous):

take the l.c.m. first and try to calculate it.

OpenStudy (anonymous):

[7x*2(8x+4)-1]/2(8x+4)=6 14x(8x+4)-1=6*2(8x+4) 112x^2+56x-1=96x+48 112x^2+56x-96x-1-48=0

OpenStudy (anonymous):

112x^2-40x-49=0

OpenStudy (anonymous):

We don't use the distributive property?

OpenStudy (anonymous):

step 1 i s wrong

OpenStudy (anonymous):

I was thinking the same but because of the 4x but I wasn't sure

OpenStudy (anonymous):

its 7x -8x-4=6 -x-4=6 -x=6+4 x=-10

OpenStudy (anonymous):

from where di 4x come?its not given in the ques.have a look

OpenStudy (anonymous):

oh the 1/2 didnt come out

OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 7x - 1/2 (8x + 4) = 6 Step 1: 7x - 4x - 2 = 6 Step 2: 3x - 2 = 6 Step 3: 3x - 2 + 2 = 6 Step 4: 3x = 6 Step 5: x = 2 Which step was the error made and why? (1 point) Select one: a. Step 1, because incorrect signs have been used b. Step 2, because the terms have been added incorrectly c. Step 3, because a number has been added only to one side of the equation d. No error was made

OpenStudy (anonymous):

I think your tight first step is wrong so the answer is A?

OpenStudy (anonymous):

first step is wrong itself

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (wolf1728):

Equation stated in the question is 7x-1/2(8x+4)=6 How does this become 7x - (8x + 4) = 6 in the answer? *************************************** Okay I see Rebella posted 2 answers and the equation is correct in the second one. Rebella, could you delete that first posting?

OpenStudy (anonymous):

Okay thank you I deleted the one with that doesn't have the 1/2

OpenStudy (wolf1728):

Thanks Rebella

OpenStudy (wolf1728):

Actually the mistake is in Step 3 Where does that +2 come from? So your answer is: c. Step 3, because a number has been added only to one side of the equation

OpenStudy (anonymous):

see...according to the eqn....first step is itself wrong.if first will be wrong..rest of the steps automatically will be wrong

OpenStudy (wolf1728):

Okay, the question states How to solve 7x-1/2(8x+4)=6 step by step Then the first 2 lines of the answer are: 7x - 1/2 (8x + 4) = 6 Step 1: 7x - 4x - 2 = 6 Where is the mistake?

OpenStudy (anonymous):

u try to solve this eqn first..can u show me first two steps?

OpenStudy (wolf1728):

The answer to the equation states 7x - 1/2 (8x + 4) = 6 Step 1: 7x - 4x - 2 = 6 Step 2: 3x - 2 = 6 Everything looks fine until Step 3

OpenStudy (anonymous):

is (8x+4) in the denominator or numerator

OpenStudy (anonymous):

it is written in a way as its in denominator.

OpenStudy (wolf1728):

The equation would be a LOT easier to solve if it were written like this: 

OpenStudy (anonymous):